OpenStudy (anonymous):

what is the length of a ramp that has a base of 96 feet and is 28 feet high ?

OpenStudy (stamp):

Draw the picture. \[Hint:\ a^2+b^2=c^2\]

OpenStudy (anonymous):

By the Pythagorean theorem, we know that for any triangle whose side lengths are 'a' and 'b', its length, c, is given by \[c = \sqrt{a ^{2}+b ^{2}}\] Thus, you get c=100.

OpenStudy (mathstudent55):